OpenStudy (anonymous):

A 1200-kg car starts from rest and accelerates with constant acceleration, traveling 200 m in 9.00 s. What is the force of the road on the car during this acceleration?

OpenStudy (sidsiddhartha):

use s=ut+1/2*a*t^2 u=0 so 200=1/2*a*t^2 at^2=400 a=400/81 now force on the car F=ma=5.925 kn